A 77.0 kg box slides down a 25.0◦
ramp with
an acceleration of 3.30 m/s2
.Find µk between the box and the ramp.
The acceleration of gravity is 9.81 m/s
2
.
1
Expert's answer
2019-12-23T11:26:59-0500



From the equation "F=ma"

"mgSin(25)-F' =ma \to(1)\newline"

Where m is the mass of the box ,F' is the frictional force acted on the box and 'a' is the acceleration

"\because F'=77.0*9.81*sin(25)-77.0*3.30 \newline
F'=65.13N"


Considering the Forces perpendicular to the ramp

"R=mgCos(25)"

solving this,

"R=77.0*9.81Cos(25) \newline
R=684.60N"


Therefore the Frictional force acting on the box,

"F'=\mu R \newline
\mu=\frac{F'}{R} \newline
\mu=0.095"